the current in the river flowed at 3 mph. that boat could travel 24 miles downstrea in 1/2 the time it took to travel 12 miles upstream, what was the speed of the boat in still water?
Answer by stanbon(75887)

the current in the river flowed at 3 mph. that boat could travel 24 miles downstrea in 1/2 the time it took to travel 12 miles upstream, what was the speed of the boat in still water?
-------------------------
Downstream DATA:
distance = 24 miles ; time = x/2 hrs ; rate = d/t = 48/x mph
----------------------
Upstream DATA:
distance = 12 miles ; time = x hrs ; rate = d/t = 12/x mph
----------------------
EQUATIONS:
Let speed of the boat in still water = "b":
b+3 = 48/x
b-3 = 12/x
---------------
Subtract 2nd from 1st to get:
6 = 36/x
x = 6 hrs
------------
Solve for "b":
b+3 = 48/6
b + 3 = 8
b = 5 mph (speed of the boat in still water)
====================
Cheers,
Stan H.